OpenStudy (anonymous):

(2-i) + (3-square root of -3)

OpenStudy (anonymous):

\[2-i + (3-\sqrt{-3})\]

OpenStudy (mertsj):

\[(2-i)+(3-\sqrt{-3})=(2-i)+(3-\sqrt{3(-1)})=2-i+3-\sqrt{3}i=5-(1-\sqrt{3})i\]

OpenStudy (mertsj):

\[5-(1-\sqrt{3})i\]